OpenStudy (anonymous):

Find a Quadratic Equation Corresponding to the Solutions Given Below. A) x=2, x=-9/2 B) x=-3+2i

OpenStudy (lgbasallote):

if x = 2 and x = -9/2 are the roots then that means x + 2 and x + 9/2 are the factors so i can rewrite this as (x+ 2)(2x+9) expand it to get your quadratic equation dpes that help?

OpenStudy (lgbasallote):

does*

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

yes, but then how i would i do B?

OpenStudy (anonymous):

think it should be \((x- 2)(2x+9)\) to give one zero at \(x=2\)

OpenStudy (anonymous):

if one zero is \(-3+2i\) the the other must be its conjugate \(-3-2i\)

OpenStudy (anonymous):

so your unenviable job is to multiply this out \[(x-(-3+2i))(x-(-3-2i))\] believe it or not, it is not nearly as hard as it looks

OpenStudy (lgbasallote):

oh yes sorry...thanks @satellite73 ...i got careless there

OpenStudy (anonymous):

in fact, it is actually very easy, you can just about do it in your head

OpenStudy (anonymous):

i can walk you through it if you like

OpenStudy (anonymous):

I'm still a little confused on how you would plug it in so it's basically just a conjugate?

OpenStudy (anonymous):

yes please walk me through it

OpenStudy (anonymous):

yes first of all if \(a+bi\) is a zero of a real polynomial, then so is \(a-bi\) the conjugate

OpenStudy (anonymous):

now we know that it must factor as \((x-(a+bi))(x-(a-bi))\) by the factor theorem if you have a zero, you can factor now it is our job to multiply this mess out but as i said it is not that hard would you like to do it with the general form \[ (x-(a+bi))(x-(a-bi))\] or the specific one \[(x-(-3+2i))(x-(-3-2i))\]?

OpenStudy (anonymous):

the specific one please

OpenStudy (anonymous):

lets to it with the specific one first don't distribute, just use "first outer inner last"

OpenStudy (anonymous):

first is obviously \(x^2\) lets hold on to the "outer inner" for a second last is \((-3+2i)(-3-2i)\) and it is always true that \((a+bi)(a-bi)=a^2+b^2\) a real number so don't think about minus signs or \(i\) or anything, it is just \[(-3+2i)(-3-2i)=3^2+2^2=13\]

OpenStudy (anonymous):

I see and i get it now. the answer is (-3+2i)(-3-2i)? because of it having a zero so then it has to have its conjugate which is a negative and plus

OpenStudy (anonymous):

now for the middle part, outer and inner you are going to get \((3-2i)x+(3+2i)x\) and when you add them the \(2i\) part goes and you are simply left with \(6x\)

OpenStudy (anonymous):

yes the zeros come in conjugate pair so the original equation must have been \[x^2+6x+13\]

OpenStudy (anonymous):

or rather \[x^2+6x+13=0\] if you solve, you will see that the zeros are \(-3\pm2i\)

OpenStudy (anonymous):

ahhh i see now. you get that conjugate pair then multiply them both together but then also with the X's to get your quadratic equation.

OpenStudy (anonymous):

thanks a lot, i get it all now.

OpenStudy (anonymous):

yes you need them both now lets do it with \(a+bi\) and \(a-bi\) so you can do it quickly if you need it on a test

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

the quadratic it came from must have been \[x^2-2ax+(a^2+b^2)=0\]

OpenStudy (anonymous):

in your example \(a=-3,b=2\) so you got \[x^2-2(-3)x+3^2+2^2=0\] or \[x^2+6x+13=0\]

OpenStudy (anonymous):

impress your teacher and do it in 5 seconds

OpenStudy (anonymous):

zeros are \(4+3i\) and \(4-3i\) equation must have been \[x^2-8x+25=0\]

OpenStudy (anonymous):

hope it is clear (more or less, may need some practice)

OpenStudy (anonymous):

do you have to foil them both once you get the roots?

OpenStudy (anonymous):

not sure what you are asking the way this worked, we were given the roots and we had to come up with the equation so yes we sort of "foiled" but we get the answer right away

OpenStudy (anonymous):

okay i see now given the roots so we could find the quadratic equation

OpenStudy (anonymous):

suppose i tell you the zeros are \(2+\sqrt{5}i\) and \(2-\sqrt{5}i\) and ask for the equation they came from no agony here, i say \[x^2-2\times 2x+2^2+\sqrt{5}^2=0\] or \[x^2-4x+9=0\]

OpenStudy (anonymous):

the reason i mentioned \(\sqrt{5}\) is because frequently you get a radical when you solve a quadratic equation

OpenStudy (anonymous):

okay i see and get everything now, thanks satellite!